OpenStudy (anonymous):

10/5√5 .... I know the answer is 2√5/5 but cannot figure out how to do it

OpenStudy (anonymous):

Multiply and divide by \(\sqrt{5}\): \[\frac{10}{5 \sqrt{5}} \times \frac{\sqrt{5}}{\sqrt{5}} \implies \frac{10 \sqrt{5}}{5 \times 5} \implies \frac{\cancel{10}^{2}\sqrt{5}}{\cancel{5} \times 5} \implies \frac{2 \sqrt{5}}{5}\]

OpenStudy (anonymous):

Thank you, I appreciate your help.

OpenStudy (anonymous):

This is called \(\color{blue}{\mathbb{Rationalazing \quad the \quad Denominator..}}\)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

Thanks I appreciate the reference as well.